OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? a. five, six, eleven b. five, nine, ten c. five, thirteen, eighteen d. five, twelve, thirteen

OpenStudy (anonymous):

try to see if they follow pythagoras theorem. see 5^2 + 6^2 = 25 + 36 = 61 which is not equal to 11^2 = 121. Try out that way and see which option is correct/

OpenStudy (anonymous):

soooo D?

OpenStudy (anonymous):

yes. :)